OpenStudy (anonymous):

For the function f(x) = 3x+2, find x such that f(x) = 14. Numerical Answers Expected!

OpenStudy (cwrw238):

solve 3x + 2 = 14

OpenStudy (anonymous):

well i dont know how?

OpenStudy (cwrw238):

the aim is first to get terms in x on one side of the = and numbers the other side so subtract 2 from each side of the above equation - what does that give?

OpenStudy (cwrw238):

3x + 2 - 2 = 14 - 2 can you simpify this?

OpenStudy (anonymous):

nope i forgot to simpify

OpenStudy (cwrw238):

whats 2 - 2? and 14 - 2 ? i'm sure you can do that

OpenStudy (anonymous):

0

OpenStudy (cwrw238):

thats 2-2 whats 14 - 2?

OpenStudy (anonymous):

12

OpenStudy (cwrw238):

ok so thats makes the equation 3x - 0 = 12 3x = 12 now divide 3x by 3 and 12 by 3 and you have x = ?

OpenStudy (cwrw238):

* 3x + 0 = 12

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what well i do now

OpenStudy (cwrw238):

x = 12/3 whats 12/3 equal to

OpenStudy (anonymous):

4

OpenStudy (cwrw238):

right thats the solution x = 4

OpenStudy (anonymous):

oh so that is the answer right

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw